Find the equilibrium concentrations of A and B for a=2 and b=2. Assume that the initial concentration of A is 1.0 MM and that no B is present at the beginning of the reaction.Kc= 3.0

Express your answers using two significant figures separated by a comma.

I made an ICE chart for this problem and I get 3.0= [2X]^2 / [1-2X]^2

I then get 3.0(1-2X)^2 = 4X^2.

Then I get 3 + 12X^2 - 12X = 4X^2.

Then 8X^2- 12X +3 = 0

Then 8X(X-1) - 4(x+ 3/4)

I get stuck after that.

Find the equilibrium concentrations of A and B for a = 1 and b = 2. Assume that the initial concentration of A is 1.0 MM and that no B is present at the beginning of the reaction. Kc= 3.0

I had trouble with this one as well.
Answer #1

I ZAT 2B Initial - Equilibrium 1-2x 2x ų = [B]² [A] ² = (2x)2 (1-2x)² 3 = 2x 1-2x 3 - 23 n = 2x 3 = x(2+253) = x 3 2+253 1 =